Question 4:In Terminal A at Sacramento airport, Peet’s has a virtual monopoly on the sale of coffee past the point of security. Inverse demand for coffee in Terminal A is given by p(q) =21−0.1q. Their cost function is given byC(q) =1q+1000 where the fixed cost is what Peet’s pays Sacramento airport in rent.( a) How many coffees will Peet’s sell? What is the price of coffee? (b) Plot the demand for coffee, mark the equilibrium price and quantity. Comment on the welfare effects of having a monopolist selling coffee in Terminal A.2 (c) Despite Peet’s having a substantial markup above its marginal cost of production, who is the ultimate recipient of these monopoly “rents” (i.e. positive economic profits)?
Answer #1

Plq) = 21-0.19 clq) = Iq +10oo. cas 2 case Th= of a monopoly: Px q = 219 - 0.19 MR = NTR = 21- 0.29 MR = MC, MC=de = 1. 21-0.